OpenStudy (anonymous):

A number, n, is added to 15 less than 3 times itself. The result is 101. Which equation can be used to find the value of n?

OpenStudy (tyler444):

options?

OpenStudy (anonymous):

3n – 15 + n = 101 3n + 15 + n = 101 3n – 15 – n = 101 3n + 15 – n = 101

OpenStudy (tyler444):

i believe the correct option is D because you start off with 3 times n then you add 15 then subtract n to get 101 hope this is helpful

OpenStudy (anonymous):

Thank You !

OpenStudy (tyler444):

your welcome

OpenStudy (anonymous):

can you help me with some more

OpenStudy (tyler444):

yes i will do my best

OpenStudy (anonymous):

What are the possible steps involved in solving the equation shown? Check all that apply. 3.5 + 1.2(6.3 – 7x) = 9.38 Add 3.5 and 1.2. Distribute 1.2 to 6.3 and –7x . Combine 6.3 and –7x . Combine 3.5 and 7.56. Subtract 11.06 from both sides. Divide both sides by 9.38.

OpenStudy (tyler444):

@micahm

OpenStudy (micahm):

A number, n, is added to 15 less than 3 times itself. The ... openstudy.com/updates/55494789e4b0324d54bb961f A number, n, is added to 15 less than 3 times itself. The result is 101. Which equation can be used to find the value of n? Delete. Cancel; Submit. Share; Tweet ... Medals 0. n + (3n - 15) = 101 or it could also be 4n - 15 = 101. Show full answer.

OpenStudy (tyler444):

I think Distribute 1.2 to 6.3 and –7x is the right option, furthermore there is another rigfht option hint: 3.5 + 1.2(6.3 −− 7x) = 9.383.5 + 7.56−8.4x = 9.3811.06−8.4x=9.38

OpenStudy (tyler444):

@kendyl513

OpenStudy (anonymous):

got it , Thank You guys

OpenStudy (tyler444):

your welcome hope we could help